The electric current in a circuit is given by I = I0 sin ( t +θ). What is the dimension  of  ? a) second   b) 1/second   c) meter / second  d) dimensionless  
2. The velocity varies with time according to the relation, v = 3t + 4. The distance travelled by the body in t = 2 s will be 
a) 10 m  b) 12 m    c) 14 m  d) 16 m  
3. When a projectile is at the highest point on its trajectory, the potential and kinetic energies are respectively 
a) maximum and minimum  b) minimum and zero  c) zero and maximum  d) maximum and zero  
4. A block of mass 2 kg starts moving when the angle of inclination of the inclined plane is 60o. If the coefficient of kinetic friction is 0.6, the frictional force is 
a) 2 N  b) 1 N    c) 4 N  d) 0.5 N  
5. Two forces F1 = (7i + 2j) N and F2 = (-5i + 3j) N act on a particle. The third force F3 that should act on the particle to make it move with constant velocity is  
a) (2i + 5j) N  b) (-2i – 5j) N   c) (–2i + 5j) N  d) (2i – 5j) N 
6. Two satellites of masses 3M and M orbit the earth in circular orbits of radii r and 3r respectively. The ratio of their speeds is 
a) 1 : 1       b) : 1     c) 3 : 1  d) 9 : 1  
 

Grade:12th pass

1 Answers

Abhishek Sharma
4 years ago
2-cv=dx/dt so x=(3t^2/2)+4t then at t=2 x=14
3-a, potential energy=mgh, kinetic energy=(1/2)mv^2, at maximum height height is max.and v less
5-b, the net force acting in x direction is 2,and in y direction is 5,so compensation for them b is suitable answer
6-a